OpenStudy (anonymous):

Which equation could be used to solve y^2+7y=18 by factoring? y(y + 7) = 18 (y + 2)(y + 9) = 0 (y + 9)(y - 2) = 0

jimthompson5910 (jim_thompson5910):

y^2+7y=18 y^2+7y-18 = 0 now find two numbers that multiply to -18 and add to 7 to help factor the left side

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

let me try y^2+7y=18 y^2+7y-18=0 (y + 9) (y - 2) = 0 so the answer is (y + 9)(y - 2) = 0

OpenStudy (anonymous):

@xo_kansasprincess_xo : I'm the kind of person who wants the people around me to be comfortable, to laugh, and to have a fabulous time - I want them to be as happy as I am :) :)